OpenStudy (anonymous):

Factor p2 - 8p - 20.

OpenStudy (anonymous):

I see youre asking about quite a few of these. Have you tried using \[x = -b \pm \sqrt{b^2 - 4ac} \over 2a \] where a b and c are \[ax^2 + bx + c\] Or in your case, a = 1, b = -8, c = -20. You get the roots \[8 \pm \sqrt{64 - 4*1*(-20)} \over 2*1\] = \[8 \pm \sqrt{144} \over 2\] which yields x = 10 and x = -2. So then you rewrite in the form (x - P)(x-Q) where P and Q are your roots. Your solution is then (x-10)(x+2) If you check you get x2 -10x + 2x -20 = x2 -8x -20